# Compilation for Dynamically Field-Programmable Qubit Arrays with Efficient and Provably Near-Optimal Scheduling

### v1

@@ V Routing: Independent Set
Not all AOD movements from one Rydberg stage to another can be performed simultaneously. The reason lies in the fundamental constraints of AOD: the order of its columns cannot change, nor can the order of rows. We consider the movements of the second stage consisting of gates $(q_{0},q_{1})$, $(q_{3},q_{8})$, $(q_{2},q_{6})$, and $(q_{4},q_{9})$, in Fig. 5a. We define a move to be a 4-tuple: $x$ and $y$ of the source, and $x$ and $y$ of the destination. Since each gate has a choice of which of its two qubits to move, there are two tuples corresponding to each gate. For example, $m_{0}=(0,1,1,2)$ and $m_{1}=(1,2,0,1)$ are both for gate $(q_{0},q_{1})$. We call them to be each other’s dual. The AOD constraints are enforced by forbidding conflicts illustrated in Fig. 5b. If the sources of two moves $m$ and $m^{\prime}$ have the same $y$, i.e., $\text{src}_{y}(m)=\text{src}_{y}(m^{\prime})$, the two qubits are picked up by the same AOD row. Then, $\text{dst}_{y}(m)=\text{dst}_{y}(m^{\prime})$ because that AOD row can only terminate at one vertical location post-movement. Similarly, if $\text{dst}_{y}(m)=\text{dst}_{y}(m^{\prime})$, then $\text{src}_{y}(m)=\text{src}_{y}(m^{\prime})$. If the qubits are picked up by different rows, their relative order must be maintained, e.g., if $\text{src}_{y}(m)>\text{src}_{y}(m^{\prime})$, then $\text{dst}_{y}(m)>\text{dst}_{y}(m^{\prime})$. In the $X$ direction, there are similar three types of conflicts.
Figure 5: Routing in Enola. a) Definition of a move as a 4-tuple. b) Conflicts between two moves. c) Compatible moves are independent sets (IS) in the conflict graph (filled vertices). After finding an IS, delete the moves and their duals from the graph. The process continues until no moves are left.
@@
These conflicts are pairwise, which means they can be encoded as edges in a graph where the vertices are the moves. We present this conflict graph in Fig. 5c. A set of compatible moves constitutes an independent set (IS) of vertices. One can utilize a maximum independent set (MIS) solver for compatible moves, but MIS is NP-hard. [^1] One can also imagine formulating the routing problem as a vertex coloring to find all compatible sets together, but this involves increasing the size of the graph and solving NP-hard coloring problems. Thus, we do not explore this possibility in this paper. In practice, we find maximal independent sets are sufficient, which can be derived by 1) putting all vertices in a list, 2) adding the first vertex to the IS, 3) removing all its neighbors from the list, and continuing 2-3). In the first box in Fig. 5c, assuming the list of vertices is sorted by indices, $m_{0}$ is added to the IS first, so its neighbors $m_{1}$, $m_{2}$, $m_{3}$, $m_{5}$, and $m_{7}$ are removed from the list. Next, $m_{4}$ is added to the IS and invalidates all the rest of vertices. So the maximal IS is $\{m_{0},m_{4}\}$ corresponding to gates $(q_{3},q_{8})$ and $(q_{2},q_{6})$. Next, $m_{0}$ and $m_{4}$, along with their duals $m_{1}$ and $m_{5}$ are deleted from the conflict graph, resulting the the second box in Fig. 5c. In the updated graph, we find the second maximal IS, $\{m_{2},m_{6}\}$. By now, all moves are deleted and the routing terminates.
@@
The runtime of maximal IS is $O(|V|+|E|)$ where $|V|$ is the number of moves, which is less or equal than the number of qubits, $n$. To construct the graph, we need to check conflicts for all pairs of vertices, which requires $O(n^{2})$ time. The longest move in each compatible set determines the AOD movement time for this set. Thus, in our compiler, the list of moves is sorted by their distance. This sorting takes $O(n\log n)$ time. Then, the maximal IS takes $O(n^{2})$ time. In summary, finding a compatible set takes $O(n^{2})$ time. In the worst case, each compatible set includes only one gate. Then, we run $O(n)$ times the procedure above until all gates in one Rydberg stage are handled, resulting in $O(n^{3})$ time. In total, there can be $O(n)$ Rydberg stages for a commutable group, so the total routing time is $O(n^{4})$. We refer to this routing approach as sortIS.
@@
To improve the scaling of sortIS, we can introduce a fixed length window when scanning the list of vertices. Instead of constructing the whole conflict graph, we only construct a graph on the first $K$ vertices in the list where $K$ is the constant window size. These vertices are the $K$ longest moves. Thus, both checking the conflicts between vertices and solving the maximal IS only take $O(K^{2})$ time. Thus, the windowed routing takes $O(n^{2}\log n+n^{2}K^{2})$ time. We refer to this routing approach as windowIS.
@@
For each compatible set of moves, the qubits need to be picked up by the AOD and dropped off to their destination interaction sites. Turning on the AOD rows and columns and ramping up the intensity for atom transfers also take time. To minimize this time, we need to consider the product structure of AOD, which is a research topic on its own [29]. Actually, we can prove that optimal routing is NP-hard from the complexity of optimizing the AOD pick-up time. In practice, we do not observe this optimality to be critical to the overall fidelity, so we defer these details to the appendix. In Enola, we apply a simple approach implemented by CodeGen in OLSQ-DPQA where the qubits are picked up row by row. The columns may shift horizontally before picking up the next row. The CodeGen just involves scanning over all the qubits to pick up, so the runtime is less than finding the compatible sets.
VI Evaluation
We implemented our proposed algorithm in Python. We employed KaMIS (v2.1) [36] for solving the maximum independent set problems. All experiments were conducted on an AMD EPYC 7V13 64-Core Processor at 2450 MHz and 128 GB of RAM. Each fidelity data point in the figures on QAOA is an average of results corresponding to 10 randomly generated graphs of the same size.
VI-A Impact of Different Settings in Enola
Fig. 6 provides the comparison of different settings in Enola on the MaxCut QAOA benchmarks. Since the scheduling is the same for all settings, the two-qubit gate fidelity term is the same. Additionally, in every setting, we use 4 atom transfers for each gate: picking up a qubit and dropping it off to the qubit it interacts with at this Rydberg stage, and the pick-up and drop-off on the way back. This means the atom transfer fidelity term is also the same for all settings. Thus, the comparison is on the decoherence term. A major improvement comes from optimizing placement, as evident by the gap between trivial placement (green triangles) and the other series. Dynamic placement (dynSA+MIS, pink cross) is slightly better than static placement (SA+MIS, blue dot). In routing, sortIS is slightly worse than MIS, as in the comparison of dynSA+sortIS (yellow star) and dynSA+MIS (pink cross). Thus, sortIS proves to be a viable replacement for MIS which is NP-hard. The windowIS method is theoretically worse than sortIS because of the limited window size. We set this size to be 1,000, larger than the scale of benchmarks in Fig. 6. In the evaluations with larger benchmarks up to 10,000 qubits, we observe a similar number of compatible move sets and an average movement distance compared to sortIS, which means windowIS is a good heuristic to speed up the compilation.
Figure 6: Decoherence fidelity term of different settings in Enola on 3-regular MaxCut QAOA circuits. ‘no SA’ means trivial placement. ‘SA’ means static placement. ‘dynSA’ means dynamic placement. ‘MIS’ means maximum independent using a solver. For these benchmarks, windowIS is the same with sortIS since the window size (1,000) is larger than the number of vertices in graph where we search for an IS.
@@ VI-A Impact of Different Settings in Enola
Figure 7: Comparison of result fidelity between Enola and OLSQ-DPQA. a) Comparison on dependency circuits. b) Comparison of total fidelity on 3-regular MaxCut QAOA circuits. c-e) Comparisons of different fidelity terms on the QAOA circuits.
@@
Figure 8: Comparison of two-qubit gate fidelity term (scattered points) on 3-regular MaxCut QAOA circuits. The total fidelity of Enola is also drawn for reference (dashes).
Figure 9: Enola runtime scaling on 3-regular MaxCut QAOA circuits. Two nearby bars correspond to the same number of qubits. The hatched bars are windowIS data (with a window size of 1,000) and the other bars are sortIS data.
VI-B Quality Comparison with Previous Works
@@
In Fig. 7, we compare Enola with OLSQ-DPQA [10] using the same set of benchmarks in that study. For dependency circuits, OLSQ-DPQA tries to execute as many gates as possible in the current front layer of the DAG, which often results in the same number of Rydberg stages as our ASAP scheduling. In some cases like the three ‘ising’ benchmarks, OLSQ-DPQA suffers from elongating the critical path because its formulation cannot explore more than one rearrangement steps between Rydberg stages, which results in a notably worse fidelity compared to Enola. In some other cases like ‘multiply_n13’ and ‘seca_n11’, it appears one rearrangement step is sufficient, so the two methods produce the same number of stages. Under this scenario, OLSQ-DPQA can potentially outperform Enola because the routing in Enola is heuristic after all and may not find the optimal compatible sets of moves.
@@
On the QAOA benchmarks, Enola clearly outperforms OLSQ-DPQA because it is able to leverage the near-optimal scheduling. We depict the comparison of overall fidelity in Fig. 7b and the three terms in Fig. 7c-e. In the two-qubit gate term, there is a significant gap between the two approaches. At 90 qubits, OLSQ-DPQA uses 14.6 stages on average whereas Enola only employs 4, a 3.7X reduction. In the atom transfer term, two approaches are similar, but Enola starts gaining advantage on larger benchmarks. It should be noted that in OLSQ-DPQA, atom transfers are not penalized in the SMT formulation. Examining its results with human eyes, there appears to be unnecessary transfers and movements. In Enola, we have explicit control over the transfers and movements. In the decoherence term, Enola is worse than OLSQ-DPQA. This is inevitable because we choose to prioritize the number of Rydberg stages, necessitating more AOD movements. Overall, Enola improves the fidelity by 5.9X compared to OLSQ-DPQA at 90 qubits.
@@
The fidelity gain of Enola is even larger when compared to heuristic methods. Q-Pilot [26] is a DPQA router that utilizes AOD only for ancilla qubits to mediate two-qubit gates between SLM qubits. It does not include a nontrivial placement solution. Atomique [27] focuses on the placement and routes the qubits with SWAP gates. In Fig. 8, we compare the two-qubit gate fidelity terms of all the approaches. Q-Pilot and Atomique result in more Rydberg stages than OLSQ-DPQA and Enola because the generation and recycling of the ancillas and the SWAPs require additional stages. At 90 qubits, Enola reduces the number of stages by 8.7X compared to Atomique and 10.5X compared to Q-Pilot. As a result, the two-qubit fidelity term of Enola (red star) is 779X higher than Atomique (green triangle) and 5806X higher than Q-Pilot (yellow cross). The total fidelity of Enola (dashes), including atom transfers and decoherence, is still higher than the two-qubit fidelity term of the two heuristics.
VI-C Runtime Scaling of Enola
@@
Fig. 9 exhibits the runtime of Enola with sortIS and windowIS on larger benchmarks, up to 10,000 qubits. Note that this is a log-log plot. Different colors inside each bar provides the portion of time spent on different tasks. The scheduling is extremely fast, invisible in the plot. For benchmarks smaller than 1,000 qubits, the runtime is dominated by the placement. Although the placement scales in $O(n)$, the constant factor is large because we would like the simulated annealing to return high-quality results. Later on, the routing portion becomes dominant due to a higher asymptotics: sortIS takes $O(n^{4})$ time and windowIS takes $O(n^{2}\log n)$ time with a constant window size. At 10,000 qubits, the sortIS approach took 1.22e4 seconds, i.e., about 3.4 hours; the windowIS approach took 1.50e3 seconds, i.e., about 25 minutes. From the data, the runtime scaling of windowIS roughly follows $O(n^{2}\log n)$: increasing the number qubits by 10X from 1,000 to 10,000, the runtime increases by 55X from 18.8 seconds to 1.04e3 seconds.
VII Related Work

### v2

@@ 5. Routing: Independent Set
In routing, we would like to parallelize the AOD movements to reduce total execution time of quantum circuits. Some movements cannot be performed simultaneously because of the fundamental constraints of AOD: the order of its columns cannot change, nor can the order of rows. We consider the movements of the second stage consisting of gates $(q_{0},q_{1})$, $(q_{3},q_{8})$, $(q_{2},q_{6})$, and $(q_{4},q_{9})$, in Fig. 5a. We define a move to be a 4-tuple: $x$ and $y$ of the source, and $x$ and $y$ of the destination. Since each gate has a choice of which one of its two qubits to move, there are two tuples corresponding to each gate. For example, $m_{0}=(0,1,1,2)$ and $m_{1}=(1,2,0,1)$ are both for gate $(q_{0},q_{1})$. We call them to be each other’s dual. The AOD constraints are enforced by forbidding conflicts illustrated in Fig. 5b. If the sources of two moves $m$ and $m^{\prime}$ have the same $y$, i.e., $\text{src}_{y}(m)=\text{src}_{y}(m^{\prime})$, the two qubits are picked up by the same AOD row. Then, $\text{dst}_{y}(m)=\text{dst}_{y}(m^{\prime})$ because that AOD row can only terminate at one vertical location post-movement. Similarly, if $\text{dst}_{y}(m)=\text{dst}_{y}(m^{\prime})$, then $\text{src}_{y}(m)=\text{src}_{y}(m^{\prime})$. If the qubits are picked up by different rows, their relative order must be maintained, e.g., if $\text{src}_{y}(m)>\text{src}_{y}(m^{\prime})$, then $\text{dst}_{y}(m)>\text{dst}_{y}(m^{\prime})$. In the $X$ direction, there are similar three types of conflicts.
@@
These conflicts are pairwise, which means they can be encoded as edges in a graph where the vertices are the moves. We present this conflict graph in Fig. 5c. A set of compatible moves constitutes an independent set (IS) of vertices. One can utilize a maximum independent set (MIS) solver for compatible moves, but MIS is NP-hard. [^1] One can also imagine formulating the routing problem as a vertex coloring on the conflict graph, where each color represents a compatible set. This formulation can find the optimal number of compatible sets whereas our approach above is greedy. However, this method involves solving NP-hard coloring problems. Furthermore, since only one of the dual moves needs to be executed, this vertex coloring problem extends beyond the conventional definition. Thus, we do not pursue this option in this work. In practice, we find maximal independent sets are sufficient, which can be derived by a simple algorithm: 1) putting all vertices in a list, 2) adding the first vertex to the IS, 3) removing all its neighbors from the list, and continuing 2-3). In the first box in Fig. 5c, assuming the list of vertices is sorted by indices, $m_{0}$ is added to the IS first, so its neighbors $m_{1}$, $m_{2}$, $m_{3}$, $m_{5}$, and $m_{7}$ are removed from the list. Next, $m_{4}$ is added to the IS and invalidates all the rest of vertices. So, the maximal IS is $\{m_{0},m_{4}\}$ corresponding to gates $(q_{3},q_{8})$ and $(q_{2},q_{6})$. Next, $m_{0}$ and $m_{4}$, along with their duals $m_{1}$ and $m_{5}$ are deleted from the conflict graph, resulting in the second box in Fig. 5c. In the updated graph, we find the second maximal IS, $\{m_{2},m_{6}\}$. By now, all moves are deleted, and the routing terminates.
Figure 5. Routing in Enola. a) Definition of a move as a 4-tuple. b) Conflicts between two moves. c) Compatible moves are independent sets (IS) in the conflict graph (filled vertices). After finding an IS, delete the moves and their duals from the graph. The process continues until no moves are left.
@@
The runtime of maximal IS is $O(|V|+|E|)$ where $|V|$ is the number of moves, which is less or equal than the number of qubits, $n$. To construct the conflict graph, we need to check conflicts for all pairs of vertices, which requires $O(n^{2})$ time. The longest move in each compatible set determines the AOD movement time for this set. Thus, in our compiler, the list of moves is sorted by their distance. This sorting takes $O(n\log n)$ time. Then, the maximal IS takes $O(n^{2})$ time. In summary, finding a compatible set takes $O(n^{2})$ time. In the worst case, each compatible set includes only one gate. Then, we run $O(n)$ times the procedure above until all gates in one Rydberg stage are handled, resulting in $O(n^{3})$ time. In total, there can be $O(n)$ Rydberg stages for a commutable group, so the total routing time is $O(n^{4})$. We refer to this routing approach as sortIS.
@@
To improve the scalability of sortIS, we can introduce a fixed length window when scanning the possible moves. Instead of constructing the whole conflict graph, we only construct a graph on the first $K$ vertices in the list where $K$ is the constant window size. These vertices are the $K$ longest moves. Thus, both checking the conflicts between vertices and solving the maximal IS only take $O(K^{2})$ time. Thus, the windowed routing takes $O(n^{2}\log n+n^{2}K^{2})$ time. We refer to this routing approach as windowIS.
@@
For each compatible set of moves, the qubits need to be picked up by the AOD and dropped off to their destination interaction sites. Turning on the AOD rows and columns and ramping up the intensity for atom transfers also take time. To minimize this time, we need to consider the product structure of AOD, which is a research topic on its own (Tan et al., 2024b). In fact, it is proven that solving DPQA routing problems optimally is NP-hard from the complexity of optimizing the AOD pick-up time, as seen in Sec. 8.5 of Ref. (Tan, 2024). In practice, we do not observe this subtask to be critical to the overall fidelity in the evaluations to follow. Thus, in Enola, we adopt a simple approach implemented in a component named CodeGen in OLSQ-DPQA where the qubits are picked up row by row. The columns may shift horizontally before picking up the next row. The CodeGen just involves scanning over all the qubits to pick up, so the runtime is less than finding the compatible sets.
@@
Figure 6. Result fidelity of Enola and OLSQ-DPQA on a) dependency circuits and b-e) 3-regular MaxCut QAOA circuits.
6. Evaluation
We implemented our proposed algorithm in Python. We employed KaMIS (v2.1) (Hespe et al., 2019) for solving the maximum independent set problems. All experiments were conducted on an AMD EPYC 7V13 64-Core Processor at 2450 MHz and 128 GB of RAM. Each fidelity data point in the figures on QAOA is an average of results corresponding to 10 randomly generated graphs of the same size.
6.1. Impact of Different Settings in Enola
Fig. 7 provides the comparison of different settings in Enola on the MaxCut QAOA benchmarks. Since the scheduling is the same for all settings, the two-qubit gate fidelity term is the same. Additionally, in every setting, we use 4 atom transfers for each gate: picking up a qubit and dropping it off to the qubit it interacts with at this Rydberg stage, and the pick-up and drop-off on the way back. This means the atom transfer fidelity term is also the same for all settings. Thus, the comparison is on the decoherence term. A major improvement comes from optimizing placement, as evident by the gap between trivial placement (green triangles) and the other series. Dynamic placement (dynSA+MIS, pink cross) is slightly better than static placement (SA+MIS, blue dot). In routing, sortIS is slightly worse than MIS, as in the comparison of dynSA+sortIS (yellow star) and dynSA+MIS (pink cross). Thus, sortIS proves to be a viable replacement for MIS which is NP-hard. The windowIS method is theoretically worse than sortIS because of the limited window size. We set this size to be 1,000, larger than the scale of benchmarks in Fig. 7. In the evaluations with larger benchmarks up to 10,000 qubits, we observe a similar number of compatible move sets, and a similar average movement distance compared to sortIS, which means windowIS is a good heuristic to speed up the compilation.
Figure 7. Decoherence fidelity term of different settings in Enola on 3-regular MaxCut QAOA circuits. ‘no SA’ means trivial placement. ‘SA’ means static placement. ‘dynSA’ means dynamic placement. ‘MIS’ means maximum independent using a solver. For these benchmarks, windowIS is the same with sortIS since the window size (1,000) is larger than the number of vertices in graph where we search for an IS.
@@ 6.1. Impact of Different Settings in Enola
Figure 8. Comparison of two-qubit gate fidelity term (scattered markers) on 3-regular MaxCut QAOA circuits. The total fidelity of Enola is also drawn for reference (dashes).
6.2. Quality Comparison with Previous Works
@@
In Fig. 6, we compare Enola with OLSQ-DPQA (Tan et al., 2024a) using the same set of benchmarks in that study. For dependency circuits [^2] Out of the 45 benchmarks in Table 1 of Ref. (Tan et al., 2024a), we display the 13 highest-fidelity circuits in Fig. 6a due to page limit. The observations we make do not change on the rest of the benchmarks., OLSQ-DPQA tries to execute as many gates as possible in the current front layer of the DAG, which often results in the same number of Rydberg stages as our ASAP scheduling. In some cases, like the three ‘ising’ benchmarks, OLSQ-DPQA suffers from elongating the critical path because its formulation cannot explore more than one rearrangement steps between Rydberg stages, which results in a notably worse fidelity compared to Enola. In some other cases like ‘multiply_n13’ and ‘seca_n11’, it happens so that one rearrangement step is sufficient, so the two methods produce the same number of stages. Under this scenario, OLSQ-DPQA can potentially outperform Enola because the routing in Enola is heuristic after all and may not find the optimal compatible sets of moves.
@@
On the QAOA benchmarks, Enola clearly outperforms OLSQ-DPQA because it is able to leverage the near-optimal scheduling. We depict the comparison of overall fidelity in Fig. 6b and the three terms in Fig. 6c-e. In the two-qubit gate term, there is a significant gap between the two approaches. At 90 qubits, OLSQ-DPQA uses 14.6 stages on average whereas Enola only employs 4, a 3.7x reduction. In the atom transfer term, two approaches are similar, but Enola starts gaining advantage on larger benchmarks. It should be noted that in OLSQ-DPQA, atom transfers are not penalized in the SMT formulation. Examining its results with human eyes, there appears to be unnecessary transfers and movements. In Enola, we only perform atom transfers and movements when necessary. In the decoherence term, Enola is worse than OLSQ-DPQA. This is inevitable because we choose to prioritize the number of Rydberg stages, necessitating more AOD movements. Overall, Enola improves the fidelity by 5.9x compared to OLSQ-DPQA at 90 qubits.
@@
The fidelity gain of Enola is even larger when compared to heuristic methods. Q-Pilot (Wang et al., 2024b) is a DPQA router that utilizes AOD only for ancilla qubits to mediate two-qubit gates between SLM qubits. It does not include a nontrivial placement solution. Atomique (Wang et al., 2024a) focuses on the placement and routes the qubits with SWAP gates. In Fig. 8, we compare the two-qubit gate fidelity terms of all the approaches. Q-Pilot and Atomique result in much more Rydberg stages than Enola because the generation and recycling of the ancillas and the SWAPs require additional stages. At 90 qubits, Enola reduces the number of stages by 8.7x compared to Atomique and 10.5x compared to Q-Pilot. As a result, the two-qubit fidelity term of Enola (red star) is 779x higher than Atomique (green triangle) and 5806x higher than Q-Pilot (yellow cross). The total fidelity of Enola (dashes), including atom transfers and decoherence, is still higher than the two-qubit fidelity term of the two heuristics. Thus, the total fidelity of the heuristics is certainly worse than Enola.
Figure 9. Enola runtime scaling on 3-regular MaxCut QAOA circuits. Two nearby bars correspond to the same number of qubits. The hatched bars are windowIS data (with a window size of 1,000) and the other bars are sortIS data.
6.3. Runtime Scaling of Enola
@@
Fig. 9 exhibits the runtime of Enola with sortIS and windowIS (default) on larger benchmarks, up to 10,000 qubits. Note that this is a log-log plot. Different colors inside each bar provides the portion of time spent on different tasks. The scheduling is extremely fast, invisible in the plot. For benchmarks smaller than 1,000 qubits, the runtime is dominated by the placement. Although the placement scales in $O(n)$, the constant factor is larger than routing. Later on, the routing portion becomes dominant due to a higher asymptotic: sortIS takes $O(n^{4})$ time and windowIS takes $O(n^{2}\log n)$ time with a constant window size. At 10,000 qubits, the sortIS approach took 1.22e4 seconds, i.e., about 3.4 hours; the windowIS approach took 1.50e3 seconds, i.e., about 25 minutes. From the data, the runtime scaling of windowIS roughly follows $O(n^{2}\log n)$: increasing the number qubits by 10x from 1,000 to 10,000, the runtime increases by 55x from 18.8 seconds to 1.04e3 seconds.
6.4. Potential Impact of Multiple AODs
As noted in Ref. (Wang et al., 2024a), it is feasible to equip DPQA with multiple AODs to facilitate the simultaneous movement of different sets of atoms. The AOD ordering constraints apply only to atoms within the same AOD, allowing multiple AODs to further enhance parallelism in the routing task. Since this multi-AOD configuration has not been experimentally demonstrated, we provide only a rough estimation of its potential impact using a simple round-robin assignment strategy. For example, if the routing stage identifies four compatible sets of moves, $M_{0}$ to $M_{3}$, and there are two AODs available, we can utilize AOD_0 to execute $M_{0}$ and AOD_1 to execute $M_{1}$, simultaneously. The duration of this ‘trunk’ of movements ($M_{0}$ and $M_{1}$), is determined by the longer of the two durations. Subsequently, as the next trunk, AOD_0 implements $M_{2}$ while AOD_1 executes $M_{3}$. The total duration of routing is then the cumulative duration of all trunks. This parallelism among multiple AODs reduces the overall computation time, leading to lower decoherence error and improved overall fidelity. On the 90-qubit QAOA benchmarks, we observe that utilizing two, three, and four AODs enhances the overall fidelity by factors of 1.54x, 1.79x, and 1.94x, respectively.
7. Related Work